A jet lands on a aircraft carrier at 55 mile (9m/s).
(a)what is its acceleration if it stops in 7 sec.
(b)what is the displacement of the plane which it is stopping.
1
Expert's answer
2020-03-04T09:58:57-0500

a)


"a=\frac{v}{t}=\frac{9}{7}=1.3\frac{m}{s^2}"


b)


"s=0.5at^2=0.5vt"

"s=0.5(9)(7)=31.5\ m"